1. If there is a 20% chance we will get a -3% return, a 30% chance of getting a 5% return, and a 50% chance of getting 15% return. What is the standard deviation for the investment?

*Hint: You should find E(R) first and calculate standard deviation accordingly.

Please round your answer to the fourth decimal. If your answer is 1.11%, then you should input "0.0111".
Answer #1

Expected return=Respective return*Respective probability

=(0.2*-3)+(0.3*5)+(0.5*15)

=8.4%

probability Return probability*(Return-Expected Return)^2
0.2 -3 0.2*(-3-8.4)^2=25.992
0.3 5 0.3*(5-8.4)^2=3.468
0.5 15 0.5*(15-8.4)^2=21.78
Total=51.24%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(51.24)^(1/2)

=0.0716(Approx)